OpenStudy (anonymous):

simplify ^3√8 please show work

OpenStudy (anonymous):

\[\sqrt[3]{8}=8^\frac{1}{3}\] \[8=4*2=2^2*2=2^3\] \[2^{3*\frac{1}{3}}=2^1=2\]